A hoist operated by an electric motor raises a load of 300kg vertically at a steady speed of 0.2ms-1. Frictional resistance can be taken to be constant at 1200N. What is the power required?



1
Expert's answer
2021-03-16T11:38:01-0400

By the definition of the power, we have:


"P=Fv=(F_g+F_{fr,r})v=((m_{hoist}+m_{load})g+F_{fr,r})v,""P=((500\ kg+300\ kg)\cdot9.8\ \dfrac{m}{s^2}+1200\ N)\cdot0.2\ \dfrac{m}{s},""P=1.81\cdot10^3\ W=1.81\ kW."